Soybean meal is 14% protein; cornmeal is 7% protein. How many pounds of each should be mixed together in to get 280-lb mixture that is 9% protein?
How many pounds of the cornmeal should be in the mixture?
__ pounds

How many pounds of the soybean meal should be in the mixture?
__ pounds

Answer by stanbon(57387)

Soybean meal is 14% protein; cornmeal is 7% protein.
How many pounds of each should be mixed together to
get 280-lb mixture that is 9% protein?
------------
Quantity Eq: S + C = 280 lb
Protein Eq: 0.14S+0.07C = 0.09*280
----------------------
Multiply thru 1st by 14
Multiply thru 2nd by 100
-----------------------
14S + 14C = 14*280
14S + 7C = 9*280
---
Subtract and solve for "C":
7C = 5*280
C = 200 lb (amt. of cornmeal needed in the mix)
---
Since S+C = 280, S = 80 lb (amt. of soybean needed in the mix)
=============================================
Cheers,
Stan H.
==============


How many pounds of the cornmeal should be in the mixture?
__ pounds
How many pounds of the soybean meal should be in the mixture?
__ pounds